You are considering adding a new software title to those published by your highly successful software company. If you add the new product it will use capacity on your disk duplicating machines that you had planned on using for your flagship product, Battlin Bobby. You had planned on ising the unused capacity to start selling Bab on the West coast in two years. You would eventually have had to purchase additional duplicating machines 10 years drom today but using the capacity for your new product will require moving this purchase up to 2 years from today. If the new machines will cost $114,000 and can be expenses under Section 179 your marginal tax rate is 21 percent and your cost of capital is 14 percent what is the opportunity cost associated with using the unised capacity the new product?
Answer #1

Opportunity cost will be the difference between the PV of the net cost of the machine.

Net cost of the machine = cost - section 179 deduction = 114,000 - 21%*114,000 = 90,060

Opportunity cost = -90,060/(1+14%)^2 + 90,060/(1+14%)^10 = -69,298.25 - 24,293.13 = -45,005.12